A farmer is going to build a rectangular pen. He has two plans to choose from. In Plan A, the length of the rectangle is four times the width. In Plan B, the length of the rectangle is 9 meters more than the width.
For what values of the width will the perimeter in Plan A be greater than that in Plan B?
Use w for the width (in meters), and solve your inequality for w .








Answer by chessace(471)

A: L = 4 W, Perim = 2 * (L + W) = 10 W
B: L = 9 + W, Perim = 2 * (2 W + 9) = 4 W + 18.
A > B: 10 W > 4 W + 18
-4W: 6 W > 18.
/6: W > 3 meters.